OpenStudy (anonymous):

Find the midpoint of the segment with endpoints (-3,2) and (3,-2)

hartnn (hartnn):

the mid-point formula is just the average of x and y co-ordinates.

OpenStudy (anonymous):

I know the formula.

hartnn (hartnn):

x= (x1+x2)/2 y = (y1+y2)/2

hartnn (hartnn):

here, x1 = -3, x2 =3 y1 = 2 y2 =-2 just plug in values!

hartnn (hartnn):

first you have x= (x1+x2)/2 so, what is (-3 + 3) /2 =.... ?

OpenStudy (anonymous):